OpenStudy (anonymous):

Can someone help me get this derivative. I did it, but I keep getting stuck.

OpenStudy (anonymous):

\[x^2+3e^x/2e^x-x\]

OpenStudy (anonymous):

I get \[4e^x-x^2-2xe^x/(2e^x-x)^2\]

ganeshie8 (ganeshie8):

\(\huge \frac{x^2+3e^x}{2e^x-x}\)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

its like that ?

ganeshie8 (ganeshie8):

ohk... you may use product rule or quotient rule. wats ur fav ? :)

OpenStudy (anonymous):

I'm using quotient

ganeshie8 (ganeshie8):

\(\huge (\frac{f}{g})' = \frac{gf' - g'f}{g^2}\)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\(\huge \frac{x^2+3e^x}{2e^x-x}\) \(\huge \frac{(2e^x-x)(x^2+3e^x)' - (2e^x-x)'(x^2+3e^x)}{(2e^x-x)^2}\)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\(\huge \frac{x^2+3e^x}{2e^x-x}\) \(\huge \frac{(2e^x-x)(x^2+3e^x)' - (2e^x-x)'(x^2+3e^x)}{(2e^x-x)^2}\) \(\huge \frac{(2e^x-x)(2x+3e^x) - (2e^x-1)(x^2+3e^x)}{(2e^x-x)^2}\)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

rest is just algebra :)

OpenStudy (anonymous):

ok thats what I got, but i get 4e^x-x^2-2xe^x/(2e^x-x)^2

OpenStudy (anonymous):

I checked on a website and it said my answer was wrong

OpenStudy (anonymous):

am i wrong?

ganeshie8 (ganeshie8):

for the top i am getting \(\large -2e^xx^2-x^2+xe^x+3e^x\)

OpenStudy (anonymous):

\[x^a * x^b = x^{a+b}\] \[\therefore e^x*e^x = e^{2x}\] i see no e^{2x} terms

OpenStudy (anonymous):

when I expand out the top I get \[4xe^x-2x^2-6e^{2x} -3xe^x-2x^2e^x-6e^2x-x^2-3e^x\]

OpenStudy (anonymous):

oh, I will try it again

ganeshie8 (ganeshie8):

e^2x cancels away right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

umm after i get \[xe^x-2x^2+6e ^{2x}-2x^2e^x-6e^{2x}+x^2+3e^x\]

ganeshie8 (ganeshie8):

looks good, combine like terms

OpenStudy (anonymous):

-2e^xx^2-x^2+xe^x+3e^x

OpenStudy (anonymous):

i get that

ganeshie8 (ganeshie8):

thats right, dont forget the denominator... :)

OpenStudy (anonymous):

cant I factor anything out ?

OpenStudy (anonymous):

when I put into this site http://www.wolframalpha.com/input/?i=dy%2Fdx+of+ln%28%28cos%28x^2%2B4%29%2Fcos%28x^2%2B9%29%29^5%29 I get a different answer

ganeshie8 (ganeshie8):

thats a different question ?

ganeshie8 (ganeshie8):

i see sin/cos/logs in ur link :o

OpenStudy (anonymous):

sorry, if you replace it with this question it would give u a different answer

OpenStudy (anonymous):

sorry i gave u the wrong link

ganeshie8 (ganeshie8):

here is the correct link :- http://www.wolframalpha.com/input/?i=d%2Fdx+%28x%5E2%2B3e%5Ex%29%2F%282e%5Ex-x%29

OpenStudy (anonymous):

yes, why do they take out the x and e?

ganeshie8 (ganeshie8):

its giving exact same answer :) just u dint enter the d/dx thing correctly in ur old link...

ganeshie8 (ganeshie8):

oly wolfram knows that -_-

ganeshie8 (ganeshie8):

its looking bit compact/neat i guess like that..

OpenStudy (anonymous):

oh ok. bc i thought my answer was wrong and i kept freaking out

ganeshie8 (ganeshie8):

ahh happens ;) simple way to feed wolfram is to use tick notation (function)'

ganeshie8 (ganeshie8):

enter below :- ( (x^2+3e^x)/(2e^x-x) )'

ganeshie8 (ganeshie8):

it interprets the derivative correctly http://www.wolframalpha.com/input/?i=%28+%28x%5E2%2B3e%5Ex%29%2F%282e%5Ex-x%29+%29%27

OpenStudy (anonymous):

oh ok. ugh

OpenStudy (anonymous):

thank you for your help.

ganeshie8 (ganeshie8):

np :)